OpenStudy (mackenziebroom15):

J and M are base angles of isosceles trapezoid JKLM. If mJ = 20x+9 and mM =14x+15, find mK.please help!

OpenStudy (jason1024):

k is 151 degrees

OpenStudy (mackenziebroom15):

Thanks :) how did you get that?

OpenStudy (anonymous):

20x+9=14x+15 20x-14x=15-9 6x=6 x=1 mj =20*1+9=29 m m=14*1+15=29 mk=180-29=151

OpenStudy (mackenziebroom15):

Thank you so much!